There were 690 documents checked. 31 documents had errors caused by operators. What percentage of this audit is operator errors?
Answer by stanbon(75887)

There were 690 documents checked. 31 documents had errors caused by operators. What percentage of this audit is operator errors?
--------------
x/100 = 31/690
x/100 = 0.0449
x = 4.49%
================
Cheers,
Stan H.